# Adversarial Score Identity Distillation: Rapidly Surpassing the Teacher in One Step

### v2

@@ Abstract
Score identity Distillation (SiD) is a data-free method that has achieved state-of-the-art performance in image generation by leveraging only a pretrained diffusion model, without requiring any training data. However, the ultimate performance of SiD is constrained by the accuracy with which the pretrained model captures the true data scores at different stages of the diffusion process. In this paper, we introduce SiDA (SiD with Adversarial Loss), which not only enhances generation quality but also improves distillation efficiency by incorporating real images and adversarial loss. SiDA utilizes the encoder from the generator’s score network as a discriminator, boosting its ability to distinguish between real images and those generated by SiD. The adversarial loss is batch-normalized within each GPU and then combined with the original SiD loss. This integration effectively incorporates the average “fakeness” per GPU batch into the pixel-based SiD loss, enabling SiDA to distill a single-step generator either from scratch or by fine-tuning an existing one. SiDA converges significantly faster than its predecessor when trained from scratch, and swiftly improves upon the original model’s performance after an initial warmup period during fine-tuning from a pre-distilled SiD generator. This one-step adversarial distillation method establishes new benchmarks in generation performance when distilling EDM diffusion models pretrained on CIFAR-10 (32x32) and ImageNet (64x64), achieving FID scores of 1.499 on CIFAR-10 unconditional, 1.396 on CIFAR-10 conditional, and 1.110 on ImageNet 64x64. It sets record-low FID scores when distilling EDM2 models trained on ImageNet (512x512), surpassing even the largest teacher model, EDM2-XXL, which had an FID of 1.81 with classifier-free guidance. Our SiDA’s results—achieved without classifier-free guidance—record FID scores of 2.156 for EDM2-XS, 1.669 for EDM2-S, 1.488 for EDM2-M, and 1.465 for EDM2-L, demonstrating significant improvements across all model sizes. Our open-source code will be integrated into the SiD codebase.
@@ 1 Introduction
The last decade has witnessed significant advancements in deep generative models, including generative adversarial networks (GANs) (Goodfellow et al., 2014; Reed et al., 2016; Karras et al., 2019), normalizing flows (Papamakarios et al., 2019), and variational auto-encoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014). Although VAEs are noted for their stability, they often produce blurred images, while GANs are acclaimed for their ability to generate photorealistic images despite challenges with training instability and generation diversity. These dynamics have driven research towards refining statistical distances to more accurately measure discrepancies between true and generated data distributions, especially in scenarios where traditional metrics fail due to non-overlapping supports in high-dimensional spaces (Arjovsky et al., 2017; Li et al., 2017; Zheng & Zhou, 2021).
@@
Departing from conventional methods, Diffusion-GAN (Wang et al., 2023b) reframes forward diffusion as a cost-effective, domain-agnostic data-augmentation strategy to tackle issues of non-overlapping distribution support. This approach introduces noise at various levels and utilizes statistical distances such as Jensen-Shannon divergence (JSD) for comparing distributions. The method promotes alignment between the true and generated data distributions at any stage during the forward diffusion process. It advocates for directly guiding the learning of a one-step generator through forward diffusion, moving away from the traditional dependence on iterative refinement-based sampling seen in reverse diffusion processes.
@@
Our approach builds on the existing fake score network and generator design of SiD, incorporating an adversarial loss that discriminates whether generated images are real or fake at various noisy stages of the forward diffusion process. The discriminator shares the weight with the encoder modules of the fake score network. To ensure compatibility with SiD’s pixel-level loss, we compute the discriminator loss at each spatial position of the latent encoder feature map and average across the batch samples within each GPU. This measure is integrated into the SiD loss for joint distillation and adversarial training and introduces no additional parameters. Our findings indicate this approach significantly improves iteration efficiency by about an order of magnitude when training models from scratch, and delivers unprecedentedly low FID scores when initialized from the SiD distilled checkpoints.
@@ 2 Related Work
Numerous techniques have been proposed to accelerate diffusion models (Zheng et al., 2022; Lyu et al., 2022; Luhman & Luhman, 2021; Salimans & Ho, 2022; Zheng et al., 2023; Meng et al., 2023; Kim et al., 2023). Recently, one-step and few-step diffusion distillation methods have gained prominence, achieving the speed of GANs while retaining the robust capabilities of diffusion models. These methods can be broadly classified into two categories: data-free (Luo et al., 2023b; Nguyen & Tran, 2024; Zhou et al., 2024b) and those requiring real images or teacher-synthesized noise-image pairs (Liu et al., 2023; Song & Dhariwal, 2023; Luo et al., 2023a; Kim et al., 2023; Sauer et al., 2023; Xu et al., 2023; Yin et al., 2024b). The latter frequently incorporate GAN-based adversarial learning as a critical component to enhance the generation performance of baseline distillation methods, which vary from consistency models (Song et al., 2023) to score distillation sampling (Poole et al., 2023; Wang et al., 2023c). Without this adversarial component, performance often degrades significantly (Kim et al., 2023; Sauer et al., 2023; Yin et al., 2024a).
@@
Diffusion-GAN introduced a novel strategy that matches the distribution between true and generated data at various stages of the forward diffusion process by introducing noise at different signal-to-noise ratios. This corrupts the observed data and aligns the distributions at multiple points throughout the diffusion pathway, using JSD for comparison to ensure a close alignment of the model distribution with the data distribution.
@@ 3 Preliminaries
Score identity Distillation (SiD): Moving beyond the challenges of computing JSD between diffused empirical data distributions, which can be unstable to optimize, and estimating the gradient of their KL divergence—known for mode-seeking behaviors when expectations are taken with respect to the model distribution—SiD employs a model-based Fisher divergence to match their scores. More specifically, let us denote the forward transition kernel used in Gaussian diffusion models by $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})=\mathcal{N}(a_{t}\boldsymbol{x}_{0},\sigma_{t}{\bf I})$, where $a_{t}\in(0,1)$, $\sigma_{t}>0$, and the signal-to-noise ratio $a_{t}^{2}/\sigma_{t}^{2}$ monotonically decreases towards zero as $t$ progresses. SiD explicitly represents the forward diffused data and model distributions as semi-implicit distributions (Yin & Zhou, 2018; Yu et al., 2023):
@@
$=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[\frac{a_{t}^{2}}{\sigma_{t}^{4}}\|f_{\phi^{*}}(\boldsymbol{x}_{t},t)-f_{\psi^{*}(\theta)}(\boldsymbol{x}_{t},t)\|_{2}^{2}\right],$ (2)
@@
To overcome the issues of non-overlapping probability support between two distributions, Diffusion-GANs (Wang et al., 2023b) propose to match the noisy distributions $p_{\mathrm{data}}(\boldsymbol{x}_{t})$ and $p_{\theta}(\boldsymbol{x}_{t})$ with a discriminator conditioned on time-step $t$:
$\min_{G}\max_{D}\mathbb{E}_{\boldsymbol{x}_{0}\sim p(\boldsymbol{x}_{0}),\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0};t)}[\log(D(\boldsymbol{x}_{t};t))]+\mathbb{E}_{\boldsymbol{z}\sim p(\boldsymbol{z}),\boldsymbol{x}_{t}^{g}\sim q(\boldsymbol{x}_{t}\,|\,G(\boldsymbol{z});t)}[\log(1-D(\boldsymbol{x}_{t}^{g};t))],$
where $t$ is sampled from an adaptive proposal distribution based on the competition between the generator and discriminator.
@@ 4 SIDA: Adversarial Score identity Distillation
$\mathcal{L}_{\theta}(\boldsymbol{x}_{t},\phi,\psi^{*}(\theta))$ $=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[\frac{a_{t}^{2}}{\sigma_{t}^{4}}\|f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi^{*}(\theta)}(\boldsymbol{x}_{t},t)\|_{2}^{2}\right],$ (4)
@@
$\min_{\psi}\hat{\mathcal{L}}_{\psi}(\boldsymbol{x}_{t},t)=\textstyle\gamma(t)\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2},$ (5)
@@
The gradient of this loss function will either directly impact the encoder part of $f_{\psi}$ or the generator, contingent upon the optimization target—whether it is the joint loss involving fake-score estimation and discrimination or the combined loss of diffusion distillation and adversarial generation. It is also important to highlight that no additional parameters have been introduced to facilitate joint score estimation and discrimination. This streamlined approach aids in maintaining model efficiency without compromising the effectiveness of the methodology.
@@
~$\tilde{\mathcal{L}}_{\theta,t,\alpha}=\omega(t)\frac{a_{t}^{2}}{\sigma_{t}^{4}}\mathcal{L}_{\theta,t,\alpha}^{(\text{sid})},$ (6)
@@
An alternative theoretical justification for this loss has recently been explored in Luo et al. (2024), demonstrating that the SiD loss can be derived under certain assumptions about the data generation mechanism, specifically involving stop-gradient conditions. SiD employs an alternating optimization strategy, updating $\psi$ according to (5) and $\theta$ according to (6), using only generator-synthesized fake data, drawn as specified in (1), for parameter updates. This approach ensures that both components are iteratively refined with outputs generated by the model itself, without relying on real data inputs.
Since $f_{\phi}$ has been adapted to serve dual purposes—not only performing denoising but also capable of outputting a 2D discriminator map for the encoder’s latent space—we are now ready to define an adversarial generation loss. This loss compels the generator to trick the discriminator map into believing its output originates from a real image at each spatial location of this latent 2D discriminator map. Consequently, the generator will be updated using both the loss specified in (6) and a GAN’s generator loss, effectively integrating these components to bridge the gap between the pretrained score $f_{\phi}$ and the true score $f_{\phi}^{*}$, and hence enhance the realism of the generated outputs.
Loss Construction that Respects the Weighting of Diffusion Distillation Loss. Building on the discussions above, we are now prepared to synthesize all components into a comprehensive loss function. We face an additional challenge: the diffusion loss used by SiD operates at the pixel level, while the GAN loss functions at the level of the latent 2D space. This discrepancy can make it challenging to balance the losses between diffusion distillation and adversarial generation. To circumvent the need for dataset-specific adjustments and to develop a universally applicable solution, we have devised a strategy that honors the existing weighting mechanisms used in training and distilling diffusion models. This approach involves performing GPU batch pooling to calculate an average “fakeness” within each GPU batch, then incorporating this average fakeness into each pixel’s loss prior to applying the pixel-specific weights. We have found this per-GPU-batch fakeness strategy to be highly effective, demonstrating robust performance across all datasets considered in this study.
@@
$\mathcal{L}_{\theta}^{(\text{adv})}=\frac{1}{|\mathcal{B}|}\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}D(\boldsymbol{x}_{t})=\frac{1}{|\mathcal{B}|W^{\prime}H^{\prime}}\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}\sum_{i=1}^{W^{\prime}}\sum_{j=1}^{H^{\prime}}D(\boldsymbol{x}_{t})[i,j].$ (7)
@@
$\mathcal{L}_{\theta}^{(\text{sida})}(\mathcal{B})=\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}\frac{\omega(t)a_{t}^{2}}{2\sigma_{t}^{4}}\left(\lambda_{\text{sid}}\mathcal{L}_{\theta,t}^{(\text{sid})}+\lambda_{\text{adv}}{C_{c}WH}\mathcal{L}_{\theta}^{(\text{adv})}\right),$ (8)
where we set $\lambda_{\text{sid}}=100$ and $\lambda_{\text{adv}}=0.01$ by default, unless specified otherwise.
For the fake score network $f_{\psi}$, whose encoder also serves as the discriminator $D$, we define the SiDA’s loss aggregated over a GPU batch as
$\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}{\gamma(t)}(\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}+{L}_{\psi}^{(\text{adv})}),$ (9)
@@
${L}_{\psi}^{(\text{adv})}=\frac{1}{2|\mathcal{B}|W^{\prime}H^{\prime}}\sum_{\boldsymbol{y}_{t},\boldsymbol{x}_{t}\in\mathcal{B}}\sum_{i^{\prime}=1}^{W^{\prime}}\sum_{j^{\prime}=1}^{H^{\prime}}\ln(D(\boldsymbol{y}_{t})[i^{\prime},j^{\prime}])+\ln(1-D(\boldsymbol{x}_{t})[i^{\prime},j^{\prime}]).$ (10)
We summarize in Algorithm 1 the implementation of SiDA, which starts training from a pretrained diffusion model, and in Algorithm 2 the SiDA algorithm that begins training from both a pretrained diffusion model and a pre-distilled SiD generator, referred to as SiD2A (SiD-initialized SiDA).

### v4

@@ Abstract
Score identity Distillation (SiD) is a data-free method that has achieved state-of-the-art performance in image generation by leveraging only a pretrained diffusion model, without requiring any training data. However, the ultimate performance of SiD is constrained by the accuracy with which the pretrained model captures the true data scores at different stages of the diffusion process. In this paper, we introduce SiDA (SiD with Adversarial Loss), which not only enhances generation quality but also improves distillation efficiency by incorporating real images and adversarial loss. SiDA utilizes the encoder from the generator’s score network as a discriminator, allowing it to distinguish between real images and those generated by SiD. The adversarial loss is batch-normalized within each GPU and then combined with the original SiD loss. This integration effectively incorporates the average “fakeness” per GPU batch into the pixel-based SiD loss, enabling SiDA to distill a single-step generator. SiDA converges significantly faster than its predecessor when distilled from scratch, and swiftly improves upon the original model’s performance during fine-tuning from a pre-distilled SiD generator. This one-step adversarial distillation method establishes new benchmarks in generation performance when distilling EDM diffusion models, achieving FID scores of 1.499 on CIFAR-10 unconditional, 1.396 on CIFAR-10 conditional, and 1.110 on ImageNet 64x64. When distilling EDM2 models trained on ImageNet 512x512, our SiDA method surpasses even the largest teacher model, EDM2-XXL, which achieved an FID of 1.81 using classifier-free guidance (CFG) and 63 generation steps. In contrast, SiDA achieves FID scores of 2.156 for size XS, 1.669 for S, 1.488 for M, 1.413 for L, 1.379 for XL, and 1.366 for XXL, all without CFG and in a single generation step. These results highlight substantial improvements across all model sizes. Our code and checkpoints are available at https://github.com/mingyuanzhou/SiD/tree/sida.
@@ 1 Introduction
The last decade has witnessed significant advancements in deep generative models, including generative adversarial networks (GANs) (Goodfellow et al., 2014; Reed et al., 2016; Karras et al., 2019), normalizing flows (Papamakarios et al., 2019), and variational auto-encoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014). Although VAEs are noted for their stability, they often produce blurred images, while GANs are acclaimed for their ability to generate photorealistic images despite challenges with training instability and generation diversity. These dynamics have driven research towards refining statistical distances to more accurately measure discrepancies between true and generated data distributions, especially in scenarios where
Figure 1: Illustration of the SiDA algorithm. The generator loss, as defined defined in (8), is determined by fake images $\boldsymbol{x}_{g}$, true-score-net-denoised fake images $f_{\phi}(\boldsymbol{x}_{t},t)$, fake-score-net-denoised fake images $f_{\psi}(\boldsymbol{x}_{t},t)$, and the fake image logits. Meanwhile, the fake score network loss, as defined in (9), is determined by $f_{\psi}(\boldsymbol{x}_{t},t)$ and the fake-score-net encoded logits of both fake and real images.
traditional metrics fail due to non-overlapping supports in high-dimensional spaces (Arjovsky et al., 2017; Li et al., 2017; Zheng & Zhou, 2021).
@@
Departing from conventional methods, Diffusion-GAN (Wang et al., 2023b) reframes forward diffusion as a cost-effective, domain-agnostic data-augmentation strategy to tackle issues of non-overlapping distribution support. This approach introduces noise at various levels and utilizes statistical distances such as Jensen–Shannon divergence (JSD) for comparing distributions. The method promotes alignment between the true and generated data distributions at any stage during the forward diffusion process. It advocates for directly guiding the learning of a one-step generator through forward diffusion, moving away from the traditional dependence on iterative refinement-based sampling seen in reverse diffusion processes.
@@
As illustrated in Figure 1, the proposed SiDA algorithm builds on the existing fake score network and generator design of SiD, incorporating an adversarial loss that discriminates whether generated images are real or fake at various noisy stages of the forward diffusion process. The discriminator shares the weight with the encoder modules of the fake score network. To ensure compatibility with SiD’s pixel-level loss, we compute the discriminator loss at each spatial position of the latent encoder feature map and average across the batch samples within each GPU. This measure is integrated into the SiD loss for joint distillation and adversarial training and introduces no additional parameters. Our findings across teacher models of various sizes and diverse datasets indicate that SiDA improves iteration efficiency by nearly an order of magnitude when training models initialized with the teacher model. Additionally, it achieves remarkably low FID scores when initialized from the SiD-distilled checkpoints, consistently outperforming the teacher models across all tested datasets and model sizes.
@@ 2 Related Work
Numerous techniques have been proposed to accelerate diffusion models (Zheng et al., 2022; Lyu et al., 2022; Luhman & Luhman, 2021; Salimans & Ho, 2022; Zheng et al., 2023; Meng et al., 2023; Kim et al., 2023). Recently, one-step and few-step diffusion distillation methods have gained prominence, achieving the speed of GANs while retaining the robust capabilities of diffusion models. These methods can be broadly classified into two categories: data-free (Luo et al., 2023b; Nguyen & Tran, 2024; Zhou et al., 2024b) and those requiring real images or teacher-synthesized noise-image pairs (Liu et al., 2023; Song & Dhariwal, 2023; Luo et al., 2023a; Kim et al., 2023; Sauer et al., 2023; Xu et al., 2023; Yin et al., 2024b). The latter frequently incorporate GAN-based adversarial learning as a critical component to enhance the generation performance of baseline distillation methods, which vary from consistency models (Song et al., 2023) to score distillation sampling (Poole et al., 2023; Wang et al., 2023c). Without the adversarial component, performance typically declines significantly (Kim et al., 2023; Sauer et al., 2023; Yin et al., 2024a), which raises concerns about the robustness and effectiveness of the underlying base models lacking adversarial learning support. We note that the term “One-Step” in this paper differs from its usage in OSGAN (Shen et al., 2021), which refers to a non-alternating GAN training technique, not a single generator pass.
@@
Diffusion-GAN introduced a novel strategy that matches the distribution between true and generated data at various stages of the forward diffusion process. This corrupts the observed data at different signal-to-noise ratios and aligns the distributions at multiple points throughout the diffusion pathway, using JSD for comparison to ensure a close alignment of the model and data distributions.
@@ 3 Preliminaries
Score identity Distillation (SiD): Moving beyond the challenges of computing JSD between diffused empirical data distributions, which can be unstable to optimize, and estimating the gradient of their KL divergence—known for mode-seeking behaviors when expectations are taken with respect to the model distribution—SiD employs a model-based Fisher divergence to match their scores. More specifically, let us denote the forward transition kernel used in Gaussian diffusion models by $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})=\mathcal{N}(a_{t}\boldsymbol{x}_{0},\sigma_{t}{\bf I})$, where $a_{t}\in(0,1)$, $\sigma_{t}>0$, and the signal-to-noise ratio $a_{t}^{2}/\sigma_{t}^{2}$ monotonically decreases towards zero as $t$ progresses. SiD explicitly represents the forward diffused data and model distributions as semi-implicit distributions (Yin & Zhou, 2018; Yu et al., 2023b):
@@
$=\textstyle\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[\frac{a_{t}^{2}}{\sigma_{t}^{4}}\|f_{\phi^{*}}(\boldsymbol{x}_{t},t)-f_{\psi^{*}(\theta)}(\boldsymbol{x}_{t},t)\|_{2}^{2}\right],$ (2)
@@
To overcome the issues of non-overlapping probability support between two distributions, Wang et al. (2023b) propose Diffusion-GANs to match the noisy distributions $p_{\mathrm{data}}(\boldsymbol{x}_{t})$ and $p_{\theta}(\boldsymbol{x}_{t})$ with a discriminator conditioned on time-step $t$, which is sampled from an adaptive proposal distribution based on the competition between the generator and discriminator. Diffusion-GANs are trained with:
$\min_{G}\max_{D}\mathbb{E}_{\boldsymbol{x}_{0}\sim p(\boldsymbol{x}_{0}),\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0};t)}[\log(D(\boldsymbol{x}_{t};t))]+\mathbb{E}_{\boldsymbol{z}\sim p(\boldsymbol{z}),\boldsymbol{x}_{t}^{g}\sim q(\boldsymbol{x}_{t}\,|\,G(\boldsymbol{z});t)}[\log(1-D(\boldsymbol{x}_{t}^{g};t))].$
@@ 4 SIDA: Adversarial Score identity Distillation
$\mathcal{L}_{\theta}(\boldsymbol{x}_{t},\phi,\psi^{*}(\theta))$ $=\textstyle\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[\frac{a_{t}^{2}}{\sigma_{t}^{4}}\|f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi^{*}(\theta)}(\boldsymbol{x}_{t},t)\|_{2}^{2}\right],$ (4)
@@
$\textstyle\min_{\psi}\hat{\mathcal{L}}_{\psi}(\boldsymbol{x}_{t},t)=\textstyle\gamma(t)\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2},$ (5)
@@
The gradient of this loss function will either directly impact the encoder part of $f_{\psi}$ or the generator, contingent upon the optimization target—whether it is the joint loss involving fake-score estimation and discrimination or the combined loss of diffusion distillation and adversarial generation. It is also important to highlight that no additional parameters have been introduced to facilitate the joint estimation of fake scores and discrimination between true and fake images. This streamlined approach aids in maintaining model efficiency without compromising the effectiveness of the methodology.
@@
~$\tilde{\mathcal{L}}_{\theta,t,\alpha}=\textstyle\omega(t)\frac{a_{t}^{2}}{\sigma_{t}^{4}}\mathcal{L}_{\theta,t,\alpha}^{(\text{sid})},$ (6)
@@
SiD employs an alternating optimization strategy, updating $\psi$ according to (5) and $\theta$ according to (6), using only generator-synthesized fake data, drawn as specified in (1), for parameter updates. This approach ensures that both components are iteratively refined with outputs generated by the model itself, without relying on real data inputs.
An alternative theoretical justification for the SiD loss presented in (6) has been recently explored by Luo et al. (2024), demonstrating that this loss, when $\alpha=1$, can be derived under certain assumptions about the data generation mechanism, specifically involving stop-gradient conditions. However, regardless of the method intended to recover the true theoretical loss that would yield the SiD gradient used in practice, it is essential to recognize that, as long as $\psi^{*}(\theta)$ or its gradient-detached version is part of the intended theoretical loss formulation—including model-based Fisher divergence—the bias introduced by replacing $\psi^{*}(\theta)$ with $\psi$ remains generally unavoidable in practice. This highlights the importance of empirical validation to ensure that the bias correction strategy in place is effective.
Since $f_{\phi}$ has been adapted to serve dual purposes—not only performing denoising but also capable of outputting a 2D discriminator map for the encoder’s latent space—we are now ready to define an adversarial generation loss. This loss compels the generator to trick the discriminator map into believing its output originates from a diffused real image at each spatial location of this latent 2D discriminator map. Consequently, the generator will be updated using both the loss given by (6) and a Diffusion-GAN’s generator loss, effectively integrating these components to bridge the gap between the pretrained score $f_{\phi}$ and the true score $f_{\phi}^{*}$, and hence enhance the realism of the generated outputs.
Loss Construction that Respects the Weighting of Diffusion Distillation Loss. Building on the discussions above, we are now prepared to synthesize all components into a coherent loss function. We face an additional challenge: the diffusion loss used by SiD operates at the pixel level, while the GAN loss functions at the level of the latent 2D space. This discrepancy can make it challenging to balance the losses between diffusion distillation and adversarial generation. To circumvent the need for dataset-specific adjustments and to develop a universally applicable solution, we have devised a strategy that honors the existing weighting mechanisms used in training and distilling diffusion models. This approach involves performing GPU batch pooling to calculate an average “fakeness” within each GPU batch, then incorporating this average fakeness into each pixel’s loss prior to applying the pixel-specific weights. We have found this per-GPU-batch fakeness strategy to be highly effective, demonstrating robust performance across all datasets considered in this study.
@@
$\textstyle\mathcal{L}_{\theta}^{(\text{adv})}=\frac{1}{|\mathcal{B}|}\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}\ln D(\boldsymbol{x}_{t})=\frac{1}{|\mathcal{B}|W^{\prime}H^{\prime}}\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}\sum_{i=1}^{W^{\prime}}\sum_{j=1}^{H^{\prime}}\ln D(\boldsymbol{x}_{t})[i,j].$ (7)
@@
$\textstyle\mathcal{L}_{\theta}^{(\text{sida})}(\mathcal{B})=\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}\frac{\omega(t)a_{t}^{2}}{2\sigma_{t}^{4}}\left(\lambda_{\text{sid}}\mathcal{L}_{\theta,t}^{(\text{sid})}+\lambda_{\text{adv},\theta}{C_{c}WH}\mathcal{L}_{\theta}^{(\text{adv})}\right),$ (8)
where we set $\lambda_{\text{sid}}=100$ and $\lambda_{\text{adv},\theta}=0.01$ by default, unless specified otherwise. We note that while $\mathcal{L}_{\theta,t}^{(\text{sid})}$ is influenced by the discrepancy between the teacher network $\phi$ and its optimal value $\phi^{*}$, $\mathcal{L}_{\theta}^{(\text{adv})}$ interacts directly with the training data, helping to mitigate this discrepancy. For the fake score network $f_{\psi}$, whose encoder also serves as the discriminator $D$, we define the SiDA’s loss aggregated over a GPU batch as
$\textstyle\sum_{\boldsymbol{x}_{t}\in\mathcal{B}}{\gamma(t)}(\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{\text{adv},\psi}}{L}_{\psi}^{(\text{adv})}),$ (9)
@@
$\textstyle{L}_{\psi}^{(\text{adv})}=\frac{1}{2|\mathcal{B}|W^{\prime}H^{\prime}}\sum_{\boldsymbol{y}_{t},\boldsymbol{x}_{t}\in\mathcal{B}}\sum_{i^{\prime}=1}^{W^{\prime}}\sum_{j^{\prime}=1}^{H^{\prime}}\ln(D(\boldsymbol{y}_{t})[i^{\prime},j^{\prime}])+\ln(1-D(\boldsymbol{x}_{t})[i^{\prime},j^{\prime}]).$ (10)
Here, $\lambda_{\text{adv},\psi}$ is a hyperparameter that we adjust to approximately balance the loss scales of both terms in the summation shown in (9). Specifically, we set $\lambda_{\text{adv},\psi}=1$ for distilling all EDM-pretrained models and $\lambda_{\text{adv},\psi}=100$ for distilling all EDM2-pretrained models.
We provide an overview of the SiDA implementation in Algorithm 1, which begins training from a pretrained EDM diffusion model. In Algorithm 2, we outline the modifications specific to EDM2, and in Algorithm 3, we present the SiDA algorithm that initiates training from both a pretrained diffusion model and a pre-distilled SiD generator, referred to as SiD2A (SiD-initialized SiDA).